OpenStudy (anonymous):

which value of theta is tan theta equal to sin theta? pi, pi/2, pi/4 or pi/6

OpenStudy (anonymous):

Since \[ \tan\theta = \frac{\sin\theta}{\cos\theta} \]You want \(\cos\theta=1\)

OpenStudy (anonymous):

still not sure how to go about getting an answer even with that equation

OpenStudy (anonymous):

You should know what \(cos\theta\) is equal to at those given angles. If you don't, then you must learn them.

OpenStudy (anonymous):

not sure what to do with the equation

OpenStudy (anonymous):

what given angles

OpenStudy (anonymous):

\[ \pi, \pi/2, \pi/4 \text{ or } \pi/6 \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is something that I just must memorize?

OpenStudy (anonymous):

Yes, I would definitely recommend memorizing them... I can derive them for you, but it is not fun.